[figure: 9.7 cm]

HEAT CONDUCTIVITY AND OXYGEN NON–STOICHIOMETRY RELATED PHONON MODES

V. Morchshakov^{a}, K.Bärner^{a}, A. Kattwinkel^{a}, A. M. Ahmed^{a}, I. El–Kassab^{a}, G. H. Rao^{b}, I. Medvedeva^{c}

aIV Physikalisches Institut, Georg–August–Universitaet Göttingen, Bunsenstrasse

11-15, D-37073 Göttingen, Germany bInstitute of Physics, Chinese Academy of Sciences, Beijing 100080, P.R. China cInstitute for Metal Physics, Ural Division of RAS, S.Kovalevskaya 18,620219 Ekaterinburg, Russia

In oxygen deficient (δ) manganites a correlation has been found between ferromagnetism and metallic resistivity ρ, just as with stoichiometric samples. A correlation of δ and the heat conductivity k has not been presented as yet. In this contribution we present in particular heat conductivity data of Nd_{0.7} Sr_{0.3} MnO_{3−δ}. Compiling the data of this and other systems we formulate the correlation. As strong variations in the absolute magnitude of k were assigned to changes in the phonon spectra and in particluar to O–related modes [1,2], we assign the found correlation to oxygen defect modes.

[1] H. Fujishiro, M. Ikebe, Physica B 263-264 (1999) 691

[2] I. El-Kassab, A.M.Ahmed, P.Mandal, K.Bärner, A.Kattwinkel, U.Sondermann, Physica B 305 (2001) p. 233

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : K.Bärner

Address for correspondence : IV Physikalisches Institut, Georg–August–Universität Göttingen, Bunsenstrasse 11-15, D-37073 Göttingen, Germany

Email address : baerner@physik4.gwdg.de